OpenStudy (help_people):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of f[g(5)]? 312 70 168 −1

OpenStudy (welshfella):

first find g(5) by replacing the x in x^2 - 1 and working it out

OpenStudy (welshfella):

can you do that?

OpenStudy (help_people):

g(5)= 5^2-1 ?

OpenStudy (help_people):

is that what you mean?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

g(5) means the value of g when x = 5

OpenStudy (help_people):

g(5) =24

OpenStudy (welshfella):

g(5) = 25 - 1 = 24 f(g(5)) = f(24) ( replacing g(5) by 24) can you work this out?

OpenStudy (welshfella):

again just relpace the x in f(x) (3x - 2) by 24 and simplify

OpenStudy (madhu.mukherjee.946):

hey its value is 70

OpenStudy (help_people):

f(x)=(3(24) -2) @welshfella

OpenStudy (help_people):

and are you sure @madhu.mukherjee.946

OpenStudy (madhu.mukherjee.946):

yeah

OpenStudy (help_people):

could you at least show your work

OpenStudy (madhu.mukherjee.946):

ok

OpenStudy (madhu.mukherjee.946):

see g(x)=x^2-1 so g(5)=25-1=24 now,f[g(5)]=3x-2=3*24-2 =72-2=70

OpenStudy (welshfella):

yes thats correct

OpenStudy (welshfella):

you multiply the 3 by 24 first then subtract the 2

OpenStudy (help_people):

ok thank you guys xoxo so the answer is 70 @madhu.mukherjee.946 @welshfella

OpenStudy (welshfella):

yw